Derek has the opportunity to buy a money machine today. The money machine will pay Derek $28,010.00 exactly 15.00 years from today. Assuming that Derek believes the appropriate discount rate is 12.00%, how much is he willing to pay for this money machine?

Answer format: Currency: Round to: 2 decimal places.
Answer #1


solution: PV=FV/(1+r)^n PV=28010/(1+0.12)^15 PV=Amount willing to pay for machine= $ FV Future value PV Present value r disco